# Electrons herald non-classical light

### v1

@@ Abstract
One sentence summary: We demonstrate free-electron generated non-classical multiphoton states in a photonic waveguide, leveraging two- and three-particle correlations produced by quantized parametric electron-photon scattering.
@@
Quantum states of light facilitate manifold applications in communication [1], computation [2, 3], and sensing [4, 5]. As a hallmark of quantum optics, photon number states have no classical counterpart and are essential in quantum metrology [6] and information processing [7, 8]. Furthermore, advanced protocols promise the generation of more exotic states, such as Schrödinger’s cat [9] and GKP states  [10] with applications in quantum computing [11, 12, 13]. Multiple approaches and platforms have been used for number-state generation, including circuit QED systems [14, 15, 16] and superconducting structures [17] in the microwave regime, as well as spontaneous parametric down-conversion [18, 19, 20, 21, 22] in the visible to near-infrared.
@@
In the search for new sources of non-classical light, free electrons present an attractive pathway. In pioneering work, electron-excited single quantum emitters, such as color centers and point defects, have already shown antibunching photon statistics in incoherent emission [23, 24, 25]. Parametric photon generation by coherent cathodoluminescence [26, 27, 28, 29, 30, 31, 32], on the other hand, promises sophisticated photonic states without involving Fermionic excitations, preserves the quantum phase between electrons and photons [33], and allows for enhancements by tunable phase matching [34, 35, 36]. In this context, photonic integrated circuits (PICs) [37, 38, 39] offer low-loss photon transport, as well as versatile mode tailoring, facilitating efficient scattering between electrons and optical modes [40]. Combining free electrons and PICs thus provides a unique platform for the excitation, control, and observation of classical and quantum optical states at widely variable wavelengths.
@@
A key element of realizing electron-driven quantum light will be to harness the correlations and entanglement of photonic and electronic states. While entanglement remains to be demonstrated, electron-photon coincidence detection [41, 42, 43, 44, 36, 45, 46] has yielded detailed information on the joint electron-photon state in time, energy, and momentum. Post-selection of time-correlated electrons and photons was recently shown to improve x-ray elemental mapping [44], as well as photonic-mode imaging and spectroscopy [36, 45]. Furthermore, a temporal fingerprint of the interaction may yield material- and excitation-specific decay times [47, 48] and improve the signal-to-noise ratio (SNR) in electron microscopy [49, 36]. Coherent cathodoluminescence combined with post-selection is further expected to enable the preparation of electron-heralded Fock [50, 51], GKP [52, 11], and other complex quantum states of light [52, 53]. However, the preparation of non-classical optical states via parametric electron-photon interactions has not yet been shown.
@@
In the experiment, electrons at 100 keV kinetic energy pass a straight waveguide. For the given waveguide dimensions, the interaction is dominated by the quasi-TM00 (transverse-magnetic) mode at wavelengths between 1300 and 1400 nm (see simulation results in Fig. 1c). Efficient electron-photon interaction requires a stable lateral positioning of the electron beam inside the evanescent tail of the mode in vacuum. To minimize charging-induced beam deflection, the conductivity of the waveguide surface is enhanced by coating the chip surface with an Indium Tin Oxide (ITO) layer (thickness $<20$ nm) (cf. Fig. 1d).
Figure 1: Correlation measurement of waveguide-coupled photons and free electrons. a A beam of electrons passes a dielectric waveguide, generating photons that are guided and detected in a fiber-based HBT setup. The electron and photon arrival times are temporally correlated. b Enlarged sketch of the interaction region: The electron passes the waveguide surface at an impact parameter of around 250 nm to interact with the evanescent vacuum field of the waveguide modes. c Simulation of the photon spectral distribution generated in a $80\,\mu$m straight waveguide by 100-keV electrons (impact parameter 250 nm). Gray vertical lines: waveguide boundaries. d False-color image of an example photonic chip (scanning electron micrograph). The chip surface is covered with ITO (brighter surface: ITO on metal, darker surface: ITO on Si3N4 or on SiO2, side walls: SiO2) e Schematic of the correlated quantized electron energy loss and photon number. The electron (green) scatters at the optical mode, generating a random number of photons (red). After detection, the signal correlation links the photon number $k$ (vertical, blue) to the electron scattering order $m$ (horizontal, green). The generation process is expected to follow a Poissonian distribution.
The induced optical state is analyzed by coincidence detection of the generating electrons and the created photons (cf. Fig. 1b). While the photons are guided through fibers to optical single-particle detectors, the electrons are dispersed in a magnetic spectrometer and recorded with an event-based detector (Timepix3 ASIC), which delivers the kinetic energy and arrival time of each electron. The readout electronics of the electron camera further provide time stamps for the arrival time on both photon detectors for electron-photon-photon temporal correlations. We expect to detect different photon numbers $k$ in each parametric electron scattering event, with the $k$ photons being temporally linked to electron energy shifts of $m\hbar\omega_{0}$ (cf. Fig. 1e). For an ideal electron-photon state (Eq. 1), the photon number and quanta of electron energy change are identical ($k=m$), with probabilities following a Poissonian distribution $c_{k}^{2}$. In this case, combining temporal particle correlation and electron energy analysis enables heralding schemes, in which the detection of an electron in a specific energy state |$\ket{E_{0}-k\hbar\omega}$ unambiguously heralds the photon number state |$\ket{k}$. In practice, finite optical transmission and inelastic scattering that does not inject light into the waveguide will render the quantized electron energy loss $m$ an upper bound $k\leq m,$ which to reach will be a technological target.
@@
We gain access to single- and two-photon generation events and characterize the optical state via an electron-heralded optical HBT detection setup shown in Figures 1a and 2a. Two single-photon detectors A and B placed behind a 50/50 beam splitter are synchronously time-stamped by the electron detector, and every electron is assigned to the closest photon on either detectors, calculating the time delay between them. The analysis of the resulting electron-photon-photon correlation counts $N_{\textrm{e,A,B}}$ is conducted as a function of the electron-energy loss $E_{\mathrm{el}}$ and the relative time delays $\tau_{\textrm{A}}=t_{\textrm{A}}-t_{\textrm{el}}$ and $\tau_{\textrm{B}}=t_{\textrm{B}}-t_{\textrm{el}}$ between the electron and photon detection (arrival times $t_{\textrm{el}}$, $t_{\textrm{A}}$ and $t_{\textrm{B}}$). An event histogram over these degrees of freedom results in a data cube, which we investigate along cuts or in projection on different axes (cf. Fig. 2b).
@@
The electron-photon correlation as a function of $\tau_{\textrm{A}}$ and $E_{\textrm{el}}$ is given in Fig. 2c. The correlation exhibits a strong coincidence peak separated from the uncorrelated, mostly unscattered electrons by approximately 0.9 eV (corresponding to a free-space photon wavelength of $\lambda=1350-1420$ nm). Coincidences are found within a temporal interval of 2.9 ns (FWHM, full-width at half-maximum), mainly defined by the electron detection jitter $\delta t_{\textrm{el}}$  [36]. Above a constant background in $\tau_{\textrm{A}}$, the temporally confined coincidence peak consists of events in which at least one electron-induced photon was generated, and both a photon and its generating electron were detected. Generally, direct detection of higher photon numbers is possible by employing photon-number resolving detectors [56, 57, 58, 59]. The detected electron-photon coincidences primarily originate from the excitation of the detectable TM00 mode. Higher-order modes are suppressed in the chip-fiber coupling, and the weakly-excited TE00 mode is further diminished due to the detector cutoff wavelength.
@@
Electron-heralded two-photon generation can be observed via a photon A - photon B correlation trace, filtered to true coincidences between photon detector A and an electron (see Methods). The arising two-dimensional coincidence histogram in Figure 2d is dominated by two photons correlated to electrons with about 1.8 eV = $2\hbar\omega$ energy loss. The photon-photon coincidence peak has a width of   400 ps (FWHM, $\delta t_{\textrm{ph}}<300$ ps photon detector jitter). Residual coincidence events in the zero-loss region and the first loss sideband stem from false coincidences due to both electron and photon loss. These false coincidences are also discussed in Figure 4a.
@@
Figure 2: Three-particle correlation for one electron and two photons. a Schematic of the detection setup with the particle arrival times $t_{\textrm{el}},t_{\textrm{A}},t_{\textrm{B}}$ on the fiber-coupled photon detectors DA, DB and the electron detector Del and electron energy $E_{\textrm{el}}$ as measured parameters. The arrival times are correlated with each other. b Parameter space for threefold events. The data is analyzed through lineouts and integrals along different axes. c Histogram of electron-photon-photon coincidence events $N_{\textrm{e,A,B}}$ as a function of $\tau_{\textrm{A}}$ and $E_{\textrm{el}}$, summed over $\tau_{\textrm{B}}$. Electron-photon scattering causes a coincidence peak at an electron energy shifted downward by around 0.9 eV. Inset: The temporal coincidence uncertainty of around 2.9 ns FWHM represents the timing precision of electron detection. d Histogram of particle triples as a function of time delay $t_{\textrm{A}}-t_{\textrm{B}}$ and electron energy $E_{\textrm{el}}$ filtered to true coincidences between a photon A and an electron. The coincidences are dominated by electrons that generated 2 photons. (Lowest count rates $<0.1\leavevmode\nobreak\ (\textrm{s}\cdot\textrm{ns}\cdot\textrm{eV})^{-1}$ are kept in lighter colors for better contrast.) Inset: The reduced timing uncertainty of 0.4 ns FWHM follows from the better time resolution in photon detection.
@@
The inelastic electron scattering at the sample is visualized in the electron energy spectra displayed in Figure 3. The uncorrelated spectrum (c.f. Fig. 3a) contains all channels of inelastic electron scattering [60], including material-specific radiative and non-radiative effects, quasi-particle excitation [61] as well as the desired coherent photon generation into optical modes. While the electron energy-loss resulting from electron scattering at a dielectric material can be described by a continuous exponential decay in energy [54], phase-matched parametric photon generation leads to the formation of discrete sidebands, downshifted by multiples $m$ of the generated photon’s energy $\hbar\omega$ and following a population distribution described by $P_{m}=|c_{m}|^{2}$ (see Equation 2 with $k=m$ and Fig. 1e). As the electron scattering channels at the material and into the optical modes are independent of each other, the measured electron spectrum can be modeled by a convolution of the spectra resulting from broadband scattering and from coherent cathodoluminescence in guided optical modes (cf. SI and Fig. 3a, red).
@@
The spectral component governing coherent photon generation is given by the yellow curve with a peak-to-peak distance between the sidebands close to 0.9 eV. The coupling constant of the coherent electron-photon interaction $g^{\textrm{EELS}}_{0}$ can be calculated from the population distribution of the individual sidebands $P_{m}$ (c.f. Fig. 3a, violet to light blue peaks). This distribution generally only holds for a single excited optical mode and constant interaction strength. However, in the experiment, an averaging of the coupling constant is observed, caused by a broadening of the impact parameter due to a nonzero convergence angle and finite beam-focus diameter (cf. Fig. 3b,c), as well as time-dependent beam shifts. Using the population ratio between the first three sidebands in the electron spectra, we obtain an average coupling constant $\langle g^{\textrm{EELS}}_{0}\rangle=0.32$ with a standard deviation of $\Delta_{g^{\textrm{EELS}}_{0}}=0.24$ (see SI for calculations).
@@
Figure 3: Multiple scattering and heralded electron spectra. a Time-averaged (not photon-correlated) electron spectrum (blue), modeled (red) by cascaded $m$-photon multiple-loss peaks (yellow) and an exponentially decaying continuum. b Sketch of the electron-beam caustic traversing the waveguide with a convergence half-angle of $\alpha=1.1$ mrad, causing some spatial averaging of the coupling constant. c Experimentally observed decay of the mean coupling strength towards larger impact parameters (distance to the waveguide). The standard deviation is given as a gray background. The peak height distributions at different $g^{\textrm{EELS}}_{0}$ are included as insets. d Energy distributions of uncorrelated (blue) and correlated electrons (red, yellow) for $k$ detected photons (normalized to the area under the curves). The correlated spectra are obtained by summing over the background-subtracted coincidence peaks shown in Figures 2c,d. e Measured optical spectrum with a central peak width of about 100 nm, or 65 meV, normalized to the maximum value.
@@
Whereas the unheralded electron spectrum includes all inelastic scattering pathways, access to the properties of the detected optical mode is enabled through electron-photon coincidences. The spectra of electrons in coincidence to one (red) or two (yellow) detected photons (cf. Fig. 3d) closely resemble copies of the uncorrelated spectrum (blue) when normalized to the area under the curve (slight deviations in the shape of the coincidence spectra may occur due to transmission losses on the heralding photons, see SI). The central wavelength of the TM00 mode, which dominates the transmitted photons, defines the relative shift of the spectra, and the phase matching bandwidth of the optical mode with 100-keV electrons of 65 meV, corresponding to 100 nm optical wavelength (cf. Fig. 3e) moderately broadens the distributions towards higher scattering orders.
@@
While the previously calculated coupling constant $g_{0}^{\textrm{EELS}}$ can include multiple inelastic scattering pathways, and thus may be very large [62], a strong coupling to a single optical mode is highly desirable. The coupling strength $g_{0}$ of electrons to the detected TM00 mode can be determined via Equation 2, taking into account optical losses (see SI for transmission measurements). The detection probability of one or two photons $P_{1}$ and $P_{2}$ is deduced from the area of the $m=1$ and $m=2$ sidebands in the electron-photon and electron-photon-photon coincidence spectra, respectively. They define the number of single- (two-)photon excitation events in which one (two) photons were detected. The calculated $g_{0}>0.2$ is somewhat lower than the value extracted from the electron spectrum, illustrating the necessity of coincident photon detection for disentangling multiple scattering pathways and, in particular, for isolating coherent photon generation.
Beyond the spectral characterization of heralded and unheralded electronic states, we quantitatively study the electron-photon-photon coincidences and their intensity correlations $g^{(2)}$.
@@
Figure 4 analyzes the relative-time histogram $N_{\textrm{e,A,B}}(\tau_{A},\tau_{B},E_{\textrm{el}})$, that counts threefold correlations at delays $\tau_{\textrm{A}}$ and $\tau_{\textrm{B}}$ with an electron energy $E_{\textrm{el}}$. Heralding a photonic state by a specific energy change is achieved by integrating over the corresponding energy windows for single ($m=1$, Fig. 4a) and two-photon energy loss ($m=2$, Fig. 4b). True coincidences of electrons with one photon detector $N_{\textrm{e,A$\cup$B}}$ (i.e., at detector DA OR DB) lead to vertical/horizontal bands at $\tau_{i}\approx 0$, while photon-photon coincidences $N_{\textrm{A$\cap$B}}$ (excitation of detectors DA AND DB) with or without a corresponding electron emerge on the diagonal line along $\tau_{A}-\tau_{B}\approx 0$. Threefold coincidences are found at the intersection of both features (for an in-detail analysis, see SI). While $m=1$ shows electron single-photon coincidences and some photon-photon coincidences not linked to an electron, the plot for $m=2$ is dominated by two photons in coincidence with a detected electron. As the three-particle coincidences are mainly confined to the $2\hbar\omega$ loss region (cf. Figures 2b, 3d), they demonstrate the generation of two-photon states via parametric electron scattering with a two-photon energy loss. The temporal widths of the coincidence features follow the respective detector jitters (cf. Fig. 2c,d). Photon-photon coincidences on the diagonal line off the center arise from photon-photon coincidences assigned to a different electron than the generating one, which may either have gone undetected or experienced higher loss.
@@
The correlation between the detected electron energy and the generated photon number renders our scheme a versatile heralded photon-number source. The intrinsic heralding efficiency quantifies how the energy-resolved electron detection predicts the number of generated photons $\eta_{\textrm{i}}^{I}=N_{\textrm{i,j}}/(N_{\textrm{j}}\eta_{\textrm{i}}^{d}T_{\textrm{i}})$ [63]. This is the probability that, given a heralding particle $j$ is detected, the heralded particle $i$ also exists. In this measurement setup, the heralding efficiency is calculated for the case of observing $k$ photons when selecting an energy window around a specific electron energy loss $m=1,2$ and vice versa ($\eta_{\textrm{i}}^{d}T_{\textrm{i}}$: transmission losses and detector efficiencies). Using the count rates from Figure 4a, the efficiencies reach $\eta^{I}_{\textrm{e}}>40\%$ for photon-heralded electrons ($m=1,2$) and $\eta^{I}_{\textrm{A$\cup$B}}=10\%$ and $\eta^{I}_{\textrm{A$\cap$B}}=0.3\%$ for electron-heralded single photons and two photons, respectively. Experimentally, $\eta^{I}_{\textrm{A$\cap$B}}$ and $\eta^{I}_{\textrm{A$\cup$B}}$ are limited by spectral overlap of the electron-loss peaks (cf. Fig. 3a), which leads to some mixing in the heralding of photon-number states. Further improvements of the heralding efficiency could, thus, include an enhanced energy resolution [64], and a better coupling ideality to the detected mode [55].
@@
Figure 4: Statistical analysis of generated photons. a,b Electron-photon-photon correlation events $N_{\textrm{e,A,B}}$ as a function of photon-electron time delays, selected by electron energy: a scattering order $m=1$, and b scattering order $m=2$. c Violation of the Cauchy-Schwartz Inequality ($\gamma>1$) for the electron-photon interaction at different time delays $\tau$ (points connected for visibility). d Unheralded photon-photon intensity correlation for varying electron current. e Electron-heralded photon-photon intensity correlation filtered to the energy region $m=1$ as a function of both time delays. Inset: Time-averaged $g^{(2)}$ for $m=1$. f Photon-photon intensity correlation as a function of the number of photon-heralding electrons $q$ between the two heralded photons without energy selection (top) and for $m=1$ (bottom).
@@
The quality of a heralded signal is further characterized by the coincidence-to-accidental ratio $\textrm{CAR}=\frac{R_{\textrm{sig}}-R_{\textrm{acc}}}{R_{\textrm{acc}}}$ [63], relating the signal $R_{\textrm{sig}}$ and background $R_{\textrm{acc}}$ rates of coincidences in a specific energy loss region $m$. A CAR of $>30$ is measured for single photons in coincidence with an electron at $m=1$, showing good fidelity in generating heralded single photons. The CAR value for electron-heralded two-photon events reaches $>150$ for $m=2$, confirming that detected two-photon states are generated predominantly by double-loss of single electrons and not by random coincidences of two single-photon generation events.

### v2

@@ Abstract
One sentence summary: We demonstrate free-electron-generated non-classical multiphoton states in a photonic waveguide, leveraging two- and three-particle correlations produced by quantized parametric electron-photon scattering.
@@
Quantum states of light facilitate manifold applications in communication [Ursin2007], computation [Kok2007, Zhong2020], and sensing [Degen2017, Ganapathy2023]. As a hallmark of quantum optics, photon number states have no classical counterpart and are essential in quantum metrology [Holland1993] and information processing [Broome2013, Wang2020]. Furthermore, advanced protocols promise the generation of more exotic states, such as Schrödinger’s cat [Ourjoumtsev2007] and GKP states [Gottesman2001] with applications in quantum computing [Baranes2023, Campagne-Ibarcq2020, Reglade2024]. Multiple approaches and platforms have been used for number-state generation, including circuit QED systems [Geremia2006, Zhou2012, Uria2020] and superconducting structures [Hofheinz2008] in the microwave regime, as well as spontaneous parametric down-conversion [Waks2006, Ourjoumtsev2006, Cooper2013, Tiedau2019, Harder2016] in the visible to near-infrared.
@@
In the search for new sources of non-classical light, free electrons present an attractive pathway. In pioneering works, electron-excited single quantum emitters, such as color centers and point defects, have already shown antibunching photon statistics in incoherent emission [Bourrellier2016, Fiedler2023, Tizei2013]. Parametric photon generation by coherent cathodoluminescence [Vesseur2007, Christopher2020, Brenny2014, Coenen2014, Coenen2017, Sannomiya2020, Scheucher2022a], on the other hand, promises sophisticated photonic states without involving Fermionic excitations, preserves the quantum phase between electrons and photons [Kfir2021], and allows for enhancements by tunable phase matching [Kfir2020, Muller2021, Feist2022]. In this context, photonic integrated circuits (PICs) [Wang2020a, Elshaari2020, Liu2021] offer low-loss photon transport, as well as versatile mode tailoring, facilitating efficient scattering between electrons and optical modes [Henke2021]. Combining free electrons and PICs thus provides a unique platform for the excitation, control, and observation of classical and quantum optical states at widely variable wavelengths.
@@
A key element of realizing electron-driven quantum light will be to harness the correlations and entanglement of photonic and electronic states. In particular, electron-photon coincidence detection [Kruit1984, Graham1986, Bendana2011, Jannis2019, Feist2022, Varkentina2022, Yanagimoto2023] has yielded detailed information on the joint electron-photon state in time, energy, and momentum. Post-selection of time-correlated electrons and photons was recently shown to improve x-ray elemental mapping [Jannis2019], as well as photonic-mode imaging and spectroscopy [Feist2022, Varkentina2022]. Furthermore, a temporal fingerprint of the interaction may yield material- and excitation-specific decay times [Varkentina2023, Taleb2024] and improve the signal-to-noise ratio (SNR) in electron microscopy [Koppell2025, Feist2022]. Coherent cathodoluminescence combined with post-selection is further expected to enable the preparation of electron-heralded Fock [Kfir2019, BenHayun2021], GKP [Dahan2023, Baranes2023], and other complex quantum states of light [Dahan2023, Sirotin2024]. However, the preparation of non-classical optical states via parametric electron-photon interactions has not yet been shown.
@@
correspond to a Poissonian process of quantized energy exchanges producing $k$ photons near a central optical frequency $\omega_{0}$, with a dimensionless coupling constant $g_{0}=\sqrt{\int d\omega|g_{\omega}|^{2}}$ obtained by integration over the spectral coupling strength density $g_{\omega}$.
For the designed structure, the interaction is dominated by the quasi-TM00 (transverse-magnetic) mode at wavelengths between 1300 and 1400 nm, while other excited optical modes are suppressed (see simulation results in Fig. 1c).
@@
[…]
@@
We gain access to single- and two-photon generation events and characterize the optical state via an electron-heralded optical HBT detection setup shown in Figures 1a and 2a. Two single-photon detectors A and B placed behind a 50/50 beam splitter are synchronously time-stamped by the electron detector, and every electron is assigned to the closest photon on either detector, calculating the time delay between them. The analysis of the resulting electron-photon-photon correlation counts $N_{\textrm{e,A,B}}$ is conducted as a function of the electron-energy loss $E_{\mathrm{el}}$ and the relative time delays $\tau_{\textrm{A}}=t_{\textrm{A}}-t_{\textrm{el}}$ and $\tau_{\textrm{B}}=t_{\textrm{B}}-t_{\textrm{el}}$ between the electron and photon detection (arrival times $t_{\textrm{el}}$, $t_{\textrm{A}}$ and $t_{\textrm{B}}$). An event histogram over these degrees of freedom results in a data cube, which we investigate along cuts or in projection on different axes (cf. Fig. 2b).
@@
The electron-photon correlation as a function of $\tau_{\textrm{A}}$ and $E_{\textrm{el}}$ is given in Fig. 2c. The correlation exhibits a strong coincidence peak separated from the uncorrelated, mostly unscattered electrons by approximately 0.9 eV (corresponding to a free-space photon wavelength of $\lambda=1350-1420$ nm). Coincidences are found within a temporal interval of 2.9 ns (FWHM, full-width at half-maximum), mainly defined by the electron detection jitter $\delta t_{\textrm{el}}$ [Feist2022]. Above a constant background in $\tau_{\textrm{A}}$, the temporally confined coincidence peak consists of events in which at least one electron-induced photon was generated, and both a photon and its generating electron were detected. The detected electron-photon coincidences primarily originate from the excitation of the detectable TM00 mode. Higher-order modes are suppressed in the chip-fiber coupling, and the weakly-excited TE00 mode is further diminished due to the detector cutoff wavelength.
@@
Electron-heralded two-photon generation can be observed via a photon A - photon B correlation trace, filtered to coincidences between photon detector A and an electron. The arising two-dimensional coincidence histogram in Figure 2d is dominated by two photons correlated to electrons with about 1.8 eV = $2\hbar\omega$ energy loss. The photon-photon coincidence peak has a width of  400 ps (FWHM, $\delta t_{\textrm{ph}}<300$ ps photon detector jitter). Residual coincidence events in the zero-loss region and the first loss sideband stem from false coincidences due to electron or photon loss.
@@
Figure 2: Three-particle correlation for one electron and two photons. a Schematic of the detection setup with the particle arrival times $t_{\textrm{el}},t_{\textrm{A}},t_{\textrm{B}}$ on the fiber-coupled photon detectors DA, DB and the electron detector Del and electron energy $E_{\textrm{el}}$ as measured parameters. The arrival times are correlated with each other. b Parameter space for threefold events. The data is analyzed through lineouts and integrals along different axes. c Bottom: Histogram of electron-photon-photon coincidence events $N_{\textrm{e,A,B}}$ as a function of $\tau_{\textrm{A}}$ and $E_{\textrm{el}}$, summed over $\tau_{\textrm{B}}$. Electron-photon scattering causes a coincidence peak at an electron energy shifted downward by around 0.9 eV. Top: The temporal coincidence uncertainty of around 2.9 ns FWHM represents the timing precision of electron detection. d Bottom: Histogram of particle triples as a function of time delay $t_{\textrm{A}}-t_{\textrm{B}}$ and electron energy $E_{\textrm{el}}$ filtered to true coincidences between a photon A and an electron. The coincidences are dominated by electrons that generated 2 photons. Top: The reduced timing uncertainty of 0.4 ns FWHM follows from the better time resolution in photon detection.
@@
The inelastic electron scattering at the sample is visualized in the electron energy spectra displayed in Figure 3. The uncorrelated spectrum (c.f. Fig. 3a) contains all channels of inelastic electron scattering [Egerton2009], including material-specific radiative and non-radiative effects, quasi-particle excitation [Schattschneider1987] as well as the desired coherent photon generation into optical modes. While the electron energy-loss resulting from electron scattering at a dielectric material can be described by a continuous exponential decay in energy [GarciadeAbajo2010], phase-matched parametric photon generation leads to the formation of discrete sidebands, downshifted by multiples $m$ of the generated photon’s energy $\hbar\omega$ and following a population distribution described by $P_{m}=|c_{m}|^{2}$ (see Equation 2 with $k=m$ and Fig. 1e). As the electron scattering channels at the material and into the optical modes are independent of each other, the measured electron spectrum can be modeled by a convolution of the spectra resulting from broadband scattering and from coherent cathodoluminescence in guided optical modes (cf. SI and Fig. 3a, red).
@@
Removing the effect of dielectric losses, the spectral component governing coherent photon generation is given by the yellow curve with a peak-to-peak distance between the sidebands close to 0.9 eV. The overall coupling constant of the electrons to the waveguide optical modes $g^{\textrm{EELS}}_{0}$ can be calculated from the population distribution of the individual sidebands $P_{m}$ (c.f. Fig. 3a, violet to light blue peaks). In the experiment, some averaging of the coupling constant is observed, caused by a broad distribution of impact parameters due to a nonzero convergence angle and finite beam-focus diameter (cf. Fig. 3b,c), as well as time-dependent beam shifts. Using the population ratio between the first three sidebands in the electron spectra, we obtain a coupling constant of $\langle g^{\textrm{EELS}}_{0}\rangle=0.32$ with a deviation of $\Delta_{g^{\textrm{EELS}}_{0}}=0.24$ (see SI for calculations).
@@
Figure 3: Multiple scattering and heralded electron spectra. a Time-averaged (not photon-correlated) electron spectrum (blue), modeled (red) by cascaded $m$-photon multiple-loss peaks (yellow) and an exponentially decaying continuum. b Sketch of the electron-beam caustic traversing the waveguide with a convergence half-angle of $\alpha=1.1$ mrad, causing some spatial averaging of the coupling constant. c Experimentally observed decay of the mean coupling strength towards larger impact parameters (distance to the waveguide). The standard deviation is given as a gray background. The peak height distributions at different $g^{\textrm{EELS}}_{0}$ are included as insets. d Energy distributions of uncorrelated (blue) and correlated electrons (red, yellow) for $k$ detected photons (normalized to the area under the curves). The correlated spectra are obtained by summing over the background-subtracted coincidence peaks shown in Figures 2c,d. e Measured optical spectrum with a central peak width of about 80 nm, or 65 meV, normalized to the maximum value.
@@
Whereas the unheralded electron spectrum includes all inelastic scattering pathways, access to the properties of the detected optical mode is enabled through electron-photon coincidences. The spectra of electrons in coincidence with one (red) or two (yellow) detected photons (cf. Fig. 3d) closely resemble copies of the uncorrelated spectrum (blue) when normalized to the area under the curve. The central wavelength of the TM00 mode, which dominates the transmitted photons, defines the relative shift of the spectra to be around 0.9 eV. A moderate spectral broadening of 65 meV per electron scattering order can be attributed to the phase-matching bandwidth of 80 nm with the optical mode (cf. Fig. 3e). This corresponds to an effective interaction length of around $40\penalty 10000\ \mu\text{m}$, in which the electron beam and the optical mode’s field can be regarded as parallel (see simulation results in the SI).
@@
While the previously calculated coupling constant $g_{0}^{\textrm{EELS}}$ includes all excitable modes, we are primarily interested in the strong coupling to a single optical mode. We gain access to its coupling constant via the intrinsic mode filtering of the optical setup and subsequent electron-photon coincidence detection. The coupling strength $g_{0}$ of electrons to the detected TM00 mode is determined taking into account optical losses (see SI for transmission measurements). The detection probabilities of one or two photons $P_{1}$ and $P_{2}$ are deduced from the area of the $m=1$ and $m=2$ sidebands in the electron-photon and electron-photon-photon coincidence spectra, respectively. They define the number of single- (two-)photon excitation events in which one (two) photons were detected. The calculated $g_{0}>0.2$ is somewhat lower than the value extracted from the electron spectrum, illustrating the necessity of coincident photon detection for disentangling multiple scattering pathways and, in particular, for isolating coherent photon generation.
Next to the spectral characterization of heralded and unheralded electronic states, we study the electron-photon-photon coincidences and their intensity correlations $g^{(2)}$. By filtering electron-photon coincidence events to electron energies of $E_{0}-m\hbar\omega$, the generated optical state is expected to contain $m$ photons, making the heralded state non-classical. The photon statistics of the measured optical state are obtained via (heralded) intensity correlation $g^{(2)}$.
@@
Figures 4a,b analyze the relative-time histogram $N_{\textrm{e,A,B}}(\tau_{A},\tau_{B},E_{\textrm{el}})$, that counts threefold correlations at delays $\tau_{\textrm{A}}$ and $\tau_{\textrm{B}}$ with an electron energy $E_{\textrm{el}}$. Heralding a photonic state by a specific energy change is achieved by integrating over the corresponding energy windows for single ($m=1$, Fig. 4a) and two-photon energy loss ($m=2$, Fig. 4b). True coincidences of electrons with one photon detector $N_{\textrm{e,A$\vee$B}}$ (i.e., at detector DA OR DB) lead to vertical/horizontal bands at $\tau_{i}\approx 0$, while photon-photon coincidences $N_{\textrm{A$\wedge$B}}$ (excitation of detectors DA AND DB) with or without a corresponding electron emerge on the diagonal line along $\tau_{A}-\tau_{B}\approx 0$. Threefold coincidences are found at the intersection of both features (for an in-detail analysis, see SI). While $m=1$ shows electron single-photon coincidences and some photon-photon coincidences not linked to an electron, the plot for $m=2$ is dominated by two photons in coincidence with a detected electron. As the three-particle coincidences are mainly confined to the $2\hbar\omega$ loss region (cf. Figures 2d, 3d), they demonstrate the generation of two-photon states via parametric electron scattering with a two-photon energy loss. The temporal widths of the coincidence features follow the respective detector jitters (cf. Fig. 2c,d). Photon-photon coincidences on the diagonal line off the center arise from photon-photon coincidences assigned to a different electron than the generating one, which may have gone undetected.
@@
The correlation between the detected electron energy and the generated photon number renders our scheme a versatile heralded photon-number source. The intrinsic heralding efficiency quantifies how the energy-resolved electron detection predicts the number of generated photons $\eta_{\textrm{i}}^{I}=N_{\textrm{i,j}}/(N_{\textrm{j}}\eta_{\textrm{i}}^{d}T_{\textrm{i}})$ [Signorini2020]. This is the probability that, given a heralding particle $j$ is detected, the heralded particle $i$ also exists. In this measurement setup, the heralding efficiency is calculated for the case of observing $k$ photons when selecting an energy window around a specific electron energy loss $m=1,2$ and vice versa ($\eta_{\textrm{i}}^{d}T_{\textrm{i}}$: transmission losses and detector efficiencies). Using the count rates from Figures 4a,b, the efficiencies reach $\eta^{I}_{\textrm{e}}>40\%$ for photon-heralded electrons ($m=1,2$) and $\eta^{I}_{\textrm{A$\vee$B}}=10\%$ and $\eta^{I}_{\textrm{A$\wedge$B}}=0.3\%$ for electron-heralded single photons and two photons, respectively. Experimentally, $\eta^{I}_{\textrm{A$\wedge$B}}$ and $\eta^{I}_{\textrm{A$\vee$B}}$ are limited by spectral overlap of the electron-loss peaks (cf. Fig. 3a), which leads to some mixing in the heralding of photon-number states. Further improvements of the heralding efficiency could, thus, include an enhanced energy resolution [Auad2022], and a better coupling ideality to the detected mode [Huang2023].
Figure 4: Statistical analysis of generated photons.
a,b
@@
Electron-photon-photon correlation events $N_{\textrm{e,A,B}}$ as a function of photon-electron time delays, selected by electron energy: a scattering order $m=1$, and b scattering order $m=2$. c Violation of the Cauchy-Schwartz Inequality ($\gamma>1$) for the electron-photon interaction at different time delays $\tau$ (points connected for visibility). d Unheralded photon-photon intensity correlation for varying electron current. e Electron-heralded photon-photon intensity correlation filtered to the energy region $m=1$ as a function of both time delays. Inset: Time-averaged $g^{(2)}$ for $m=1$. f Photon-photon intensity correlation as a function of the number of photon-heralding electrons $q$ between the two heralded photons without energy selection (top) and for $m=1$ (bottom).
@@
The quality of a heralded signal is further characterized by the coincidence-to-accidental ratio $\textrm{CAR}=\frac{R_{\textrm{sig}}-R_{\textrm{acc}}}{R_{\textrm{acc}}}$ [Signorini2020], relating the signal $R_{\textrm{sig}}$ and background $R_{\textrm{acc}}$ rates of coincidences in a specific energy loss region $m$. A CAR of >30>30 is measured for single photons in coincidence with an electron at $m=1$, showing good fidelity in generating heralded single photons. The CAR value for electron-heralded two-photon events reaches >150>150 for $m=2$, confirming that detected two-photon states are generated predominantly by double-loss of single electrons and not by random coincidences of two single-photon generation events.